OpenStudy (hellokitty17):

Identify the vertex for the graph of y = 2x2 + 8x - 3. (2, 21) (2, 17) (-2, -11) (-2, -27)

Mehek (mehek14):

\(\dfrac{-b}{2a}\) <--- formula

Mehek (mehek14):

\(\dfrac{-8}{2*2}=\dfrac{-8}{4}=-2\\x=-2\)

Mehek (mehek14):

plug that into the equation \(y=2*-2^2+8*-2-3\\2*4-16-3\\8-19=-11\)

Mehek (mehek14):

so C

OpenStudy (hellokitty17):

dang it i got it wrong in my head

oregonduck (oregonduck):

thxs

Mehek (mehek14):

did you try the formula?

Mehek (mehek14):

@OregonDuck why thanks?

OpenStudy (hellokitty17):

thanks for the help

oregonduck (oregonduck):

for the answer

Mehek (mehek14):

but it's @HelloKitty17 's question?

oregonduck (oregonduck):

yeah i am helping her she is my sis

Mehek (mehek14):

oooooooooooohhhhhhhhhh

OpenStudy (hellokitty17):

hi

Mehek (mehek14):

i didn't know that